OpenStudy (anonymous):

A curve has the slope 2x+3 at each point (x,y) on the curve. What is an equation for this curve if it passes through the point (1,2)

OpenStudy (anonymous):

I keep getting 2x^2+x-1 as my asnwer but the answer is y=x^2+3x-2

OpenStudy (anonymous):

i used y-y1 = m(x-x1) and subsitited m with 2x+3 and the x and y with (1,2).

OpenStudy (anonymous):

im not sure what im doing wrong

OpenStudy (anonymous):

write down differentiatial eq with initial condition y'=2x+3 y(1)=2 y=x^2+3x+C 2=1^2+3+C 2=4+C C=-2 y= x^2+3x-2

OpenStudy (anonymous):

to go from y' to y , I integrated it

OpenStudy (anonymous):

oh ok. I see. I was doing something else completely different. This makes sense. Thanks. It was so easy and i cant belive i couldnt figure it out